# Linear Combination of Saved Checkpoints Makes Consistency and Diffusion Models Better

### v2

@@ Abstract
Diffusion Models (DM) and Consistency Models (CM) are two types of popular generative models with good generation quality on various tasks. When training DM and CM, intermediate weight checkpoints are not fully utilized and only the last converged checkpoint is used. In this work, we find that high-quality model weights often lie in a basin which cannot be reached by SGD but can be obtained by proper checkpoint averaging. Based on these observations, we propose LCSC, a simple but effective and efficient method to enhance the performance of DM and CM, by combining checkpoints along the training trajectory with coefficients deduced from evolutionary search. We demonstrate the value of LCSC through two use cases: (a) Reducing training cost. With LCSC, we only need to train DM/CM with fewer number of iterations and/or lower batch sizes to obtain comparable sample quality with the fully trained model. For example, LCSC achieves considerable training speedups for CM (23$\times$ on CIFAR-10 and 15$\times$ on ImageNet-64). (b) Enhancing pre-trained models. Assuming full training is already done, LCSC can further improve the generation quality or speed of the final converged models. For example, LCSC achieves better performance using 1 number of function evaluation (NFE) than the base model with 2 NFE on consistency distillation, and decreases the NFE of DM from 15 to 9 while maintaining the generation quality on CIFAR-10. Our code is available at https://github.com/imagination-research/LCSC.
Keywords:
Consistency Model Diffusion Model Weight Averaging Evolutionary Search
@@ 1 Introduction
Diffusion Models (DMs) [36, 8, 40] are a newly emerged paradigm of generative modeling and have gained widespread attention rapidly in the last several years. DM achieves surprising performance in various tasks like image generation [8, 4, 33], video generation [9, 2] and 3D generation [31, 20]. However, DM requires an iterative denoising process in the generation process, which could be slow. Consistency Models (CMs) [39] are proposed to handle this dilemma. CM provides a better generation quality under one or few-step generation scenarios and is applied in many practical cases  [25, 26, 43].
@@
In this paper, we investigate the training process of DM and CM. We find that checkpoints—the model weights periodically saved during the training process—have under-exploited potential in boosting the performance of DM and CM. Specifically, there are numerous high-quality basins located near any given point along the optimization trajectory in the metric landscape. These high-quality basins cannot be reliably reached through Stochastic Gradient Descent (SGD), including its advanced variants like Adam [16]. However, we find that an appropriate linear combination of different checkpoints can locate these basins.
@@
Inspired by the observation, we propose to find a Linear Combination of Saved Checkpoints (LCSC) that optimizes the model performance. Specifically, for a set of checkpoints saved untill any point of the training process, LCSC searches for the optimal linear combination weights that optimize certain metrics (e.g., FID [7]) using an evolutionary algorithm. Note that the widely used Exponential Moving Average (EMA) [42] in DM and CM can be viewed as a special case of LCSC whose combination coefficients are determined heuristically and we will show in our experiments that EMA coefficients are sub-optimal. LCSC optimizes only in a low-dimension space and needs no back-propagation, thus could be faster than training. LCSC can optimize objectives or metrics whose gradients are hard to compute (e.g., FID), under which gradient-based optimization is not applicable. From a broader perspective, LCSC is a general method and may be used in other tasks and other neural networks.
We argue that LCSC is beneficial in all steps in the DM/CM production stage. For example, LCSC can be used to:
Reduce training cost. The training process of DM and CM is very costly. For example, training an EDM model [15] on CIFAR-10 costs about 2 days with 8 V100 GPUs, and training a CM model [39] costs about one week with 8 A100 GPUs. For datasets with higher resolution, the training cost is even larger. On ImageNet-64, SOTA DMs and CMs [4, 15, 39] are trained with a batch size of several thousand on 32 or 64 A100 GPUs for about one week, consuming very huge computational resources. With LCSC applied at the end, we can train CM/DM with many fewer iterations or smaller batch sizes and reach similar generation quality with the fully trained model, thereby reducing the computational cost of training.
@@
Enhance pre-trained models. Assuming the full training is already done, LCSC can still be applied to get a model that is better than any model in the training process. For model developers with saved checkpoints, LCSC can be directly applied. For users who can only access the final released checkpoint, they can fine-tune the model for a few more iterations and apply LCSC on these checkpoints. As DM/CM provides a flexible trade-off between generation quality and the number of generation steps, the enhanced model from LCSC could lead to either better generation quality or faster generation.
In conclusion, the contributions of this paper are as follows:
In Sec. 3, we analyze the metric landscape of DM and CM. We demonstrate that there are basins with better generation quality that cannot be reached through traditional SGD but can be achieved by linear combinations of checkpoints. We further demonstrate the sub-optimality of the commonly used EMA, illustrating the importance of exploring better ways of weight combination.
In Sec. 4, we propose our simple but effective method LCSC to address the problem. LCSC employs evolutionary search to optimize linear combination coefficients of previous checkpoints. LCSC has two use cases: (1) reducing training cost and (2) enhancing pre-trained models. LCSC needs no back-propagation and has the potential to be applied to other tasks and models.
In Sec. 5, we validate the effectiveness of LCSC on DM and CM for the two use cases. LCSC accelerates the training process of Consistency Distillation (CD) and Consistency Training (CT) by 23$\times$ and 7$\times$ on CIFAR-10, and 15$\times$ and 10.4$\times$ on ImageNet-64. Additionally, we apply LCSC to enhance the pre-trained models. For DM, LCSC by combining pre-trained checkpoints decreases the NFE from 15 to 9 while keeping the sample quality. For CM, LCSC achieves better performance with 1 NFE compared to the vanilla final model with 2 NFE. With the same NFE, models from LCSC have much better sample quality.
In Sec. 6, we present and discuss the searched patterns of combination coefficients to inspire further research on weights combination. Then we discuss the reasons why LCSC works due to the unique properties of DM and CM.
@@
Figure 2: A schematic diagram of LCSC. Given a set of checkpoints from training, LCSC use evolutionary search to find the optimal linear combination weights. LCSC can be applied on checkpoints from a training process with fewer training iterations or batch sizes and still gets similar performance (abbreviated as “Perf.”), thus reducing training cost and enabling faster training. LCSC can also enhance the final model in terms of generation quality or speed.
@@ 2 Background and Related Work
2.1 Diffusion Probabilistic Model
Let us denote the data distribution by $p_{data}$ and consider a diffusion process that perturbs $p_{data}$ with a stochastic differential equation (SDE) [40]:

### v3

@@ Abstract
Diffusion Models (DM) and Consistency Models (CM) are two types of popular generative models with good generation quality on various tasks. When training DM and CM, intermediate weight checkpoints are not fully utilized and only the last converged checkpoint is used. In this work, we find proper checkpoint merging can significantly improve the training convergence and final performance. Specifically, we propose LCSC, a simple, effective, and efficient method to enhance the performance of DM and CM, by combining checkpoints along the training trajectory with coefficients deduced from evolutionary search. We demonstrate the value of LCSC through two use cases: (a) Reducing training cost. With LCSC, we only need to train DM/CM with fewer number of iterations and/or lower batch sizes to obtain comparable sample quality with the fully trained model. For example, LCSC achieves considerable training speedups for CM (23$\times$ on CIFAR-10 and 15$\times$ on ImageNet-64). (b) Enhancing pre-trained models. When full training is already done, LCSC can further improve the generation quality or efficiency of the final converged models. For example, LCSC achieves better FID using 1 number of function evaluation (NFE) than the base model with 2 NFE on consistency distillation, and decreases the NFE of DM from 15 to 9 while maintaining the generation quality. Applying LCSC to large text-to-image models, we also observe clearly enhanced generation quality.
@@ 1 Introduction
Diffusion Models (DMs) (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2020b) as a generative modeling paradigm have rapidly gained widespread attention in the last several years, showing excellent performance in various tasks like image generation (Ho et al., 2020; Dhariwal & Nichol, 2021; Rombach et al., 2022), video generation (Ho et al., 2022; Blattmann et al., 2023), and 3D generation (Poole et al., 2022; Lin et al., 2023). DM requires an iterative denoising process in the generation process, which could be slow. Consistency Models (CMs) (Song et al., 2023) are proposed to handle this dilemma, which provides a better generation quality under one or few-step generation scenarios and is also broadly applied  (Luo et al., 2023a; b; Wang et al., 2023).
@@
In this paper, we investigate the training process of DM and CM. We find that checkpoints—the model weights periodically saved during the training process—have under-exploited potential in boosting the performance of DM and CM. In particular, within the metric landscape, we observe numerous high-quality basins located near any given point along the optimization trajectory. These high-quality basins cannot be reliably reached through Stochastic Gradient Descent (SGD), including its advanced variants like Adam (Kingma & Ba, 2014). However, we find that an appropriate linear combination of different checkpoints can locate these basins.
(a) Vanilla training
(b) LCSC
(c) Vanilla training
(d) LCSC
Figure 2: Comparison between the images generated by LCSC and vanilla training (on LCM-LoRA model (Luo et al., 2023a; b)). The prompts for the left and right images are “A car that seems to be parked illegally behind a legally parked car” and “Two women waiting at a bench next to a street” respectively. The vanilla method produces images with unrealistic front wheels (left) and an unnatural sitting posture (right), whereas LCSC produces more realistic images.
@@
Inspired by the observation, we propose Linear Combination of Saved Checkpoints (LCSC). In particular, for a set of checkpoints saved until any point of the training process, LCSC searches for the optimal linear combination coefficients that optimize certain metrics (e.g., FID (Heusel et al., 2017)) using an evolutionary algorithm. LCSC optimizes only in a low-dimension space and needs no back-propagation, thus could be faster than training. Additionally, LCSC can optimize objectives whose gradients are hard to compute, under which gradient-based optimization is not applicable. We note that the widely used Exponential Moving Average (EMA) (Szegedy et al., 2016) in DM and CM can be viewed as a linear combination method whose combination coefficients are determined heuristically. We will theoretically prove and empirically demonstrate that EMA coefficients are sub-optimal and are worse than LCSC. While this paper focuses on DM and CM, from a broader perspective, LCSC is a general method and may be expanded to other tasks and neural networks.
@@
We argue that LCSC is beneficial in all steps in the DM/CM production stage and can be used to: (a) Reduce training cost. The training process of DM and CM is very costly. On ImageNet-64, SOTA DMs and CMs (Dhariwal & Nichol, 2021; Karras et al., 2022; Song et al., 2023) require more than 10K GPU hours on Nvidia A100. For higher resolution tasks, the resource demand soars dramatically again: Stable Diffusion (Rombach et al., 2022) costs more than 150K GPU hours (approx. 17 years) on Nvidia A100. By applying LCSC at the end, we can train CM/DM with significantly fewer iterations or smaller batch sizes and reach similar generation quality with the fully trained model, thereby reducing the computational cost of training. (b) Enhance model performance. If the full training is already done, LCSC can still be applied to get a model that is better than any model in the training process. For model developers with saved checkpoints, LCSC can be directly applied. For users who can only access the final released checkpoint, they can fine-tune the checkpoint for a few more iterations and apply LCSC on these checkpoints. As DM/CM provides a flexible trade-off between generation quality and the number of generation steps, the enhanced model from LCSC could lead to either better generation quality or faster generation.
The reminder of this paper is organized as follows: In Sec. 2, we introduce the background and related works. In Sec. 3, we visualize the metric landscape of DM and CM, demonstrating the potential of checkpoint merging. Additionally, we provide theoretical analyses showing that EMA is suboptimal, and that a more flexible method for setting the merging coefficients is preferable. In Sec. 4, we present our efficient method LCSC, which flexibly adjusts the merging coefficients to find the best model through evolutionary search. A schematic diagram of LCSC is shown in Fig. 3
In Sec. 5, we conduct experiments to validate the effectiveness of LCSC on both DM and CM. The results show that LCSC can accelerate the training process of CM by up to 23$\times$. Moreover, LCSC can decrease the NFE of DM from 15 to 9 while keeping the sample quality. We highlight some of the results in Fig. 1. Additionally, we demonstrate that LCSC can be applied to text-to-image models to further improve generation quality, with some images presented in Fig. 2. In Sec. 6, we discuss the searched patterns of coefficients to inspire further research on checkpoints merging.
@@
Figure 3: A schematic diagram of LCSC. Given a set of checkpoints from training (left), LCSC use evolutionary search to find the optimal linear combination coefficients (middle). LCSC can be applied on checkpoints from a training process with fewer training iterations or batch sizes and still gets similar performance (abbreviated as “Perf.”), thus reducing training cost and enabling faster training. LCSC can also enhance the final model in terms of generation quality or speed (right).
@@ 2 Background and Related Work
In this section, we briefly introduce the foundational concepts and related works necessary for understanding this paper. A more comprehensive discussion is provided in App. B
Diffusion Probabilistic Model. Let us denote the data distribution by $p_{data}$, diffusion models (Sohl-Dickstein et al., 2015; Song et al., 2020a; Ho et al., 2020; Nichol & Dhariwal, 2021; Song et al., 2020b) learn a process that perturbs $p_{data}$ with a stochastic differential equation: